a pool is to be 30 feet wide and 90 feet long. A walkway is to be built around the entire pool. if the combined area cannot exceed 4000 square feet what are the possible widths of the walkway?

Answer by mananth(11019)

Area of pool = 30* 90 = 2700 ft^2 ( rectangle)
Area of pool + walkway = 4000 ft^2
area of walkway = 1300 ft^2
let the width of walk way be x
(30+2x)(90+2x)=4000
2700+60x+180x+4x^2=4000
4x^2+240x+2700=4000
/4
x^2+60x+675=1000
x^2+60x-325=0

x^2+65x-5x-325=0
x(x+65)-5(x+650=0
(x+65)(x-5)=0
x= 5 m
The width has to be less than 5 m